OpenStudy (anonymous):

In four rounds of a card game, you get scores of -3, 9, 8 and -3. What is your final score?

OpenStudy (jlastino):

Just add them all -3+9+8-3 = 11

OpenStudy (anonymous):

11 points.

OpenStudy (anonymous):

It says to translate to mathematical equation

OpenStudy (anonymous):

Average = 2.75

OpenStudy (jlastino):

Just assign values for each number let's say x1 x2 x3 x4 x1=-3 x2=9 x3=8 x4=-3 so x1+x2+x3+x4= Final Score

OpenStudy (anonymous):

if I add them I get 23

OpenStudy (anonymous):

2.75

OpenStudy (anonymous):

how do we get 2.75

OpenStudy (anonymous):

By finding the average. Do you know how to find the avg.?

OpenStudy (anonymous):

yeh

OpenStudy (hihi67):

jlamphere: -3+-3 = -6. -6+8 = 2. 2 + 9 is 11. now, there were 4 numbers that we used to come up with 11. we do 11/4. the total number divided by how many number it took you to get that final number. 11/4 = 2.75

OpenStudy (hihi67):

to put this into a math equation. you simply would do : (-3)+(-3)+9+8=11

OpenStudy (anonymous):

got it thank you

OpenStudy (hihi67):

no problem